eBook A seven-year medical research study reported that women whose mothers took the drug DES during pregnancy were twice as likely to develop tissue abnormalities that might lead to cancer as were women whose mothers did not take the drug a. This study compared two populations. What were the populations? - Select your answer b. Do you suppose the data were obtalned In a survey or In an experiment? Select your answer * c.For the population of women whose mothers took the drug DES during pregnancy, a sample of 3980 women showed that 63 developed tissue abnormalities that might lead to cancer. Provide a descriptive statistic (to 1 declmal) that could be used to estimate the number of women out of 1000 in this population who have tissue abnormalities. out of 1,000 women developed tissue abnormalities d. For the population of women whose mothers did not take the drug DES during pregnancy, what is the estimate (to 1 decimal) of the number of women out of 1000 who would be expected to have tissue abnormalities? (Hint: remember that women whose mothers took the drug were twice as likely to develop tissue abnormalities.) out of 1,000 women developed tissue abnormalities e. True or false: Medical studles often use a relatively large sample (In thls case, 3980) because disease occurrences can be rare and difficult to observe when only Isolated populatlons are consldered Select your answer- Check My Work Icon Koy
Answer #1

ANSWER:

a)

1) population whose mother took the drug DES during pregnancy

2) population of woman who did not take the drug.

b) experiment as the researchers were thinking about it and so decide to collect the data.

c) (no of women who developed abnormalities / sample population) * 10 = (63 / 3980) * 10 = 0.158 or 15.8%

d) % of women who have abnormalities / divided by 2 = 0.158 / 2 = 0.079 or 7.9%

e) true (medical and historical fact)